OpenStudy (anonymous):

solving square root equations: square root of x-1+3=x? help please

OpenStudy (anonymous):

Are you sure this is the equation? It's not possible for a number plus two to equal itself.

OpenStudy (anonymous):

\[\sqrt{X-1 +3=x}\]

OpenStudy (anonymous):

It's like that but with the 3 and the x out the square root

OpenStudy (anonymous):

In that case, first you want to isolate the square root, so subtract 3 from both sides\[\sqrt{x-1}+3-3=x-3\] leaving you with \[\sqrt{x-1}=x-3\] next, square both sides to get rid of the square root sign (remember, a square root is the same thing as taking something to the one-half power. Refer to your exponent rules if you're still confused).\[\sqrt{x-1}^{2}=(x-3)^{2}\] now you have \[x-1=(x-3)^{2}\] expand the right hand side of the equation using the foil method, and you get \[x-1=x ^{2}-6x+9\] subtract x-1 from both sides, leaving the left side zero. Now you have \[0=x ^{2}-7x+10\] Use the formula \[x=(-b+or -\sqrt{b ^{2}-4ac})\div2a\] to get your answer.